# STAMP-2.5D: Structural and Thermal Aware Methodology for Placement in 2.5D Integration

### v1

@@ V Conclusion and future work
In this paper, we introduced an automated floorplanning methodology that reduces thermally-induced structural stress in chiplet packages while optimizing wirelength with minimal temperature impact. Our results emphasize the importance of uniform temperature gradients for structural reliability. As future work, we aim to develop a multi-fidelity optimization framework to reduce runtime without compromising accuracy, and incorporate layer-specific CTE mismatch analysis within the metal stack. We also plan to explore interconnect technologies to further inform design strategies for robust chiplet integration.

### v2

@@ V Conclusion and future work
In this paper, we introduced an automated placement methodology that reduces thermally-induced structural stress in chiplet packages while optimizing wirelength with minimal temperature impact. Our results emphasize the importance of uniform temperature gradients for structural reliability. The methodology presented here establishes a foundational framework suitable for extensive future explorations, both component-specific and in the broader context of structural reliability within heterogeneous chiplet architectures. In the current STAMP-2.5D setup, high-fidelity ANSYS evaluation at every iteration represents the primary computational bottleneck. As future work, we aim to develop a multi-fidelity optimization framework to reduce runtime without compromising accuracy. In conclusion, we provide a foundational methodology for design space exploration. Additional constraints (signal integrity, power integrity, cost) can be seamlessly incorporated to build a unified, multi-objective optimization framework.